A person having a mass of 60 kilograms exerts a horizontal force of 200 newtons in pushing a 90-kilogram object a distance of 6 meters along a horizontal floor. He does this at constant velocity in 3 seconds.The force of friction is
1
Expert's answer
2020-12-14T07:27:28-0500

First find the acceleration

"u=0m\/s"

"s=ut+1\/2at^{2}"

"6=1/2\times a\times 3^{2}=9a/2"

"a=6\times 2/9=4/3 ms^{-2}"

"F=ma=90\times 4/3=120N"

The force of friction "=200N-120N=80N"